### v3

@@ 1 Introduction
However, it is well known Cappelli:1987xt ; Gepner:ADE , that the total list of $N=2$ superconformal minimal models includes also the models with pairing of type D and E. The problem we consider in this paper is the explicit construction of fields for the orbifolds of products that contain minimal models of type D and E as factors.
@@
For the second pair of mirror models, the first admissible group is the one used by D. Gepner when constructing the 3-generation model. We find the dual group and construct the mirror orbifold. For both models we write out $(a,c)$ fields and compare them also with the De Rham cohomology of the corresponding Calabi-Yau manifolds. At the end of the section we briefly discuss the further quotient of the second orbifold of these pair by the cyclic permutation group acting on the three $\textbf{E}_{7}$ factors to obtain $(c,c)$ and $(a,c)$ fields in Gepner’s three-generation model.
@@ 2 Models of $N=(2,2)$ SCFT and ADE classification
The fields of the model can be distributed over certain products of minimal representations of the holomorphic and anti-holomorphic $N=2$ Virasoro superalgebra. For a given $k$ there exist finitely many representations of such kind. They are generated from primary states $\Phi_{q}^{l}$ with $l=0,\dots,k$ and $q=-l,-l+2,\dots,l$, which are the highest vectors of $N=2$ Virasoro superalgebra. In $NS$ sector the primary state $\Phi_{q}^{l}$ has a conformal dimension and $U(1)$ charge:
@@
and in $R$-sector the corresponding conformal weight and charge are
@@
Recall that chiral-primary and anti-chiral-primary states
@@
appear in $NS$-sector when $q=\pm l$, so that the corresponding representations become shorter. Composing the holomorphic and anti-holomorphic chiral or anti-chiral primary states above we can build in general so-called $(c,c)$, $(a,c)$, $(c,a)$ and $(a,a)$ sets of fields which are endowed with ring structures and have important geometric interpretation Lerche:1989uy .
In the models of $N=(2,2)$ SCFT, the primary fields are composed of the appropriately paired holomorphic and anti-holomorphic primary states. In the modular bootstrap approach the particular way of pairing can be read off from the partition function. As it was found by D.Gepner Gepner:ADE the most general modular invariant partition function takes the form
$Z(\tau)=\sum_{l,\bar{l}=0}^{k}\sum_{\begin{subarray}{c}q,\bar{q}=-(k+1);\\ l+q+s=0,\ \bar{l}+\bar{q}+\bar{s}=0\ \text{mod}\ 2\end{subarray}}^{k+2}\sum_{\begin{subarray}{c}s,\bar{s}=0\end{subarray}}^{3}N_{l,\bar{l}}M_{q,\bar{q}}L_{s,\bar{s}}\chi_{q,s}^{l}(\tau)\bar{\chi}_{\bar{q},\bar{s}}^{\bar{l}}(\bar{\tau}),$ (5)
where $\chi_{q,s}^{l}$ is the character of the highest weight representation $\mathcal{H}^{l}_{q}$ of $N=2$ Virasoro superalgebra, which is calculated over the subspace with a fixed fermion number $s$. The matrices $N_{l,\bar{l}},M_{q,\bar{q}},L_{s,\bar{s}}$ provide modular invariance of the partition function (see also Gepner:Parafermion ).
@@
which includes the simplest diagonal invariant of type $\textbf{A}_{k+1}$ for all integer levels $k\geq 1$. For even $k$, two series of off-diagonal invariants of type D arise: $\textbf{D}_{2j+2}$ for $k=4j$, and $\textbf{D}_{2j+1}$ for k=4j2k=4j-2. Besides these infinite series, there are also three exceptional cases: $\textbf{E}_{6},\textbf{E}_{7},$ and $\textbf{E}_{8}$ for $k=10,16,$ and 28 respectively.
In (5) the parameter $q$ takes through all possible values $q=-(k+1),\dots,(k+2)$ and depends on $l$ through the parity condition: l+q=0mod 2l+q=0\ $\text${mod}\ 2. However, the minimal representations of the $N=2$ Super-Virasoro algebra are defined only for $q=-l,-l+2,\dots,l$. To clarify the interpretation of the remaining terms in the partition function, we can rely on the isomorphism of $N=2$ Virasoro superalgebra representations established in Feigin
@@
It allows us to identify the states
$\Phi_{q}^{l}\simeq\Phi_{q+(k+2)}^{k-l},\quad\Phi_{q}^{l}\simeq\Phi_{q+2(k+2)}^{l},$ (8)
Below in our article, especially in section 5 with examples, the states $\Phi^{l}_{q}$ with $|q|>l$ may arise. They should be understood as being identified with $\Phi^{k-l}_{q\pm(k+2)}$. Moreover, we restrict for simplicity ourselves to the case of diagonal matrices $M$ and $L$.
@@
Of course the holomorphic factorization we depicted in (9) should not be taken literally. The summation of quantum group indices, which ensures mutual locality of fields within correlation functions, is certainly implied and will be implied in subsequent similar expressions. But these details are not very important for the discussion in the paper. So, in what follows we will use, rather symbolic expressions like (9) instead of explicit formulas.
Taking into account the identification (8) we can assume that the charge $q$ runs:
$q=-(k+1),\dots,(k+2),\quad q+l=0\ (\text{mod}\ 2).$ (10)
@@ 2.1 Example: spectrum of $\textbf{E}_{7}$ minimal model
Let us demonstrate the spectrum of the minimal model $\textbf{E}_{7}$ with $k=16$ as an example. The partition function of the corresponding $\widehat{sl}(2)$ WZW model is
$Z_{E_{7}}=\left|\chi_{0}+\chi_{16}\right|^{2}+\left|\chi_{4}+\chi_{12}\right|^{2}+\left|\chi_{6}+\chi_{10}\right|^{2}+\left|\chi_{8}\right|^{2}+\left(\chi_{2}+\chi_{14}\right)\bar{\chi}_{8}+\chi_{8}\left(\bar{\chi}_{2}+\bar{\chi}_{14}\right)$ (11)
The set of $NS$ primary fields is shown in the table 1.
The first column contains the term from the partition function, which corresponds to the fields in the second column. Note that some fields have left and right states with different charges: $\bar{q}=q\pm(k+2)$. It arises from the identification rules (8). It is not difficult to write out the primary fields of the $R$-sector as well.
@@ 3.1 Group of symmetry of the composite model
$G_{\text{tot}}=\prod_{i=1}^{r}G_{i}=\left\{\prod_{i}\hat{g}_{i}^{w_{i}},w_{i}\in\mathbb{Z},\hat{g}_{i}=\exp\left(i2\pi J_{(i),0}\right)\right\}$ (15)
The $G_{\text{tot}}$ depends on the modular invariant (6) in each copy of $M_{k_{i}}$. In the $\textbf{D}_{2j+2}$ type series for $k=4j$, as well as in $\textbf{E}_{7}$ and $\textbf{E}_{8}$ the parameters $l_{i},\bar{l}_{i}$ run through only even values. Then the charges $q_{i}$ of primary state are also only even. It follows that the corresponding element $w_{i}\in G_{i}$ is defined modulo $(k_{i}+2)/2$, then $G_{i}\simeq\mathbb{Z}_{(k_{i}+2)/2}$. For modular invariants of type $\textbf{A}_{k_{i}+1}$ and $\textbf{D}_{2j+1}$ with $k_{i}=4j-2$ as well as for the exceptional $\textbf{E}_{6}$, there are no restrictions on the parity of the charges. Then the corresponding group is $G_{i}\simeq\mathbb{Z}_{(k_{i}+2)}$. The case of diagonal invariants BBP has a group $G_{\text{tot}}=\prod_{i=1}^{r}\mathbb{Z}_{k_{i}+2}$
Below in the text, it will be useful to regard the elements of the group $w_{i}\in\mathbb{Z}_{(k_{i}+2)/2}$ as even elements in $\mathbb{Z}_{k_{i}+2}$. We will explain this restriction in more detail in section 4.
The admissible group is defined as follows
$G_{\text{adm}}=\left\{\vec{w}=(w_{1},\dots,w_{r})\bigg{|}\quad w_{i}\in G_{i},\quad\sum_{i}^{r}\frac{w_{i}}{k_{i}+2}\in\mathbb{Z}\right\}.$ (16)
As will be seen, the constraint
@@ 3.2 Spectral Flow twisting and mutual locality of fields
In order to implement the first step we use the spectral flow realization of primary states BP ; BBP ; IIAB . It goes as follows.
@@
For values $t\in\mathbb{Z}+\frac{1}{2}$ spectral flow maps $NS$ and $R$ sectors into each other, and for $t\in\mathbb{Z}$ it acts inside $NS$ or $R$ sectors.
The extension of the spectral flow automorphism above to the $N=2$ Virasoro superalgebra representations Lerche:1989uy generates so-called twisted representations Feigin . In particular, the initial highest weight vector $\Phi^{l}_{q}$ becomes the twisted ones after the spectral flow operator $U^{t}$ action. At the same time, exactly the characters of these twisted representations appear in modular invariant partition function of Gepner. It raises the question of whether we can construct the standard highest weight vector using the states of the twisted representations?
@@
gives spectral flow realization of the $NS$ primary state $\Phi^{l}_{q}$ of the individual minimal model, where q=l2tq=l-2t. For the purposes of the orbifold construction, we need to extend this formula. Namely, the state
@@
gives the spectral flow realization of the $NS$ primary state $\Phi^{\tilde{l}}_{\tilde{q}}$, where $\tilde{l}=k-l$, $\tilde{q}=k+2+l-2t$. It is clear that the $R$ primary state is given by $U^{\frac{1}{2}}$ twisting of the $NS$ primary fields above. The validity of formulas (19), (20) follows from the structure of singular vectors of unitary, integrable representations of the $N=2$ super-Virasoro algebra Feigin . Below we generalize these formulas to construct the primary fields in the orbifold model.
Notice that for $t=k+2$ we obtain the state which can be identified with the original chiral primary one $\Phi^{l}_{c}$. However, it is implemented as a twisted one. From this point of view, spectral flow has a periodicity condition
$U^{k+2}\simeq 1,$ (21)
which means that the parameter $t$ is defined modulo $k+2$.
In terms of spectral flow realization of the primary states, the field $\Phi_{q}^{l}(z)\bar{\Phi}^{\bar{l}}_{q}(\bar{z})$ from the minimal model $M_{k}$ with explicit ADE invariant can be realized as follows
@@
Now we use the spectral flow realizations (19), (20), (22) to build twisted states from the states of the composite model for each element of the admissible group $\vec{w}\in G_{\text{adm}}$. In $NS$ sector the twisted primary field is given by
@@
$\bar{V}_{\bar{t}}^{\vec{\bar{l}}}(\bar{z})=\prod_{i}^{r}\bar{V}_{\bar{t}_{i}}^{\bar{l}_{i}}(\bar{z}),\quad\bar{V}_{\bar{t}_{i}}^{\bar{l}_{i}}(z)=\left(UG_{-\frac{1}{2}}^{-}\right)_{i}^{\bar{t}_{i}}\bar{\Phi}^{\bar{l}_{i}}_{c}(z),\quad 0\leq\bar{t}_{i}\leq\bar{l}_{i}$ (24)
@@
Then the primary states of the minimal model factors of the composite model can be represented as
@@
where the $\hat{\Phi}_{q_{i}}^{l_{i}}(z)$ is uncharged with respect to current $J_{i}(z)$. The complete primary field of the model $M_{k_{i}}$ is an appropriate product of the factor above and an antiholomorphic factor that can be represented similarly (as we have already mentioned the appropriate product must include quantum group indices summation).
@@
$\exp\left[i\sum_{i}\left(\sqrt{\frac{k_{i}}{k_{i}+2}}w_{1i}+\frac{q_{1i}}{\sqrt{k_{i}\left(k_{i}+2\right)}}\right)\phi_{i}(z)\right]\times\\ \times\exp\left[i\sum_{j}\left(\sqrt{\frac{k_{j}}{k_{j}+2}}w_{2j}+\frac{q_{2j}}{\sqrt{k_{j}\left(k_{j}+2\right)}}\right)\phi_{j}(0)\right].$ (31)
Applying the well-known result for exponents of free fields:
@@
This equation is completely analogous to the equation obtained for composite models of type A in BBP ; Parkhomenko:2022kju . Thus, the off-diagonality of the fields (23) with respect to the indices $\vec{l},\vec{\bar{l}}$, arising for D and E invariants, affects the mutual locality condition only through the allowed set of charges $q_{i}$.
@@
Thus, by solving the mutual locality equations (33) with respect to $q_{i}$ and taking into account descendants, we can explicitly determine all fields of the orbifold model. It is easy to see that in this way we have obtained a set of (quasi-)local fields that also satisfy the other axioms of the conformal bootstrap. In other words, we obtain a model of a $N=(2,2)$ superconformal field theory.

### v4

@@ 1 Introduction
However, it is well known Cappelli:1987xt ; Gepner:1988 , that the total list of $N=2$ superconformal minimal models includes also the models with pairing of type D and E. The problem we consider in this paper is the explicit construction of fields for the orbifolds of products that contain minimal models of type D and E as factors.
@@
For the second pair of mirror models, the first admissible group is the one used by D. Gepner when constructing the 3-generation model. We find the dual group and construct the mirror orbifold. For both models we write out $(a,c)$ fields and compare them also with the De Rham cohomology of the corresponding Calabi-Yau manifolds. At the end of the section we briefly discuss the farther quotient of the second orbifold of these pair by the cyclic permutation group acting on the three $\textbf{E}_{7}$ factors to obtain $(c,c)$ and $(a,c)$ fields in Gepner’s three-generation model.
@@ 2 Models of $N=(2,2)$ SCFT and ADE classification
The fields of the model can be distributed over certain products of minimal representations of the holomorphic and anti-holomorphic $N=2$ Virasoro superalgebra. For a given $k$ there exist finitely many representations of such kind. They are generated from primary states $\Phi_{q}^{l}$ with $l=0,\dots,k$ and $q=-l,-l+2,\dots,l$, which are the highest vectors of $N=2$ Virasoro superalgebra. In the $NS$ sector, the primary state is determined by the conditions
$\begin{gathered}L_{n}\Phi^{l}_{q}=0,\ J_{n}\Phi^{l}_{q}=0,\ n\geq 1,\quad G_{r}^{\pm}\Phi^{l}_{q}=0,\ r\geq\frac{1}{2}\\ L_{0}\Phi^{l}_{q}=\Delta_{l,q}\Phi^{l}_{q},\ J_{0}\Phi^{l}_{q}=Q_{l,q}\Phi^{l}_{q}.\end{gathered}$ (2)
In the $R$ sector, the definition of the primary state is distinguished by the action of the fermionic generators of the $N=2$ Virasoro algebra:
$G_{n}^{\pm}\Phi^{l}_{q}=0,\ n\geq 1,\quad G_{0}^{+}\Phi^{l}_{q}=0.$ (3)
The $NS$ sector primary state $\Phi_{q}^{l}$ has a conformal dimension and $U(1)$ charge:
@@
and in the $R$ sector the corresponding conformal weight and charge are
@@
A key class of representations in the $NS$ sector are those generated by chiral-primary states, which are defined by an additional constraint:
$G^{+}_{-1/2}\Phi^{l}_{q}=0.$ (6)
Similarly, an anti-chiral primary state in the $NS$ sector must satisfy, in addition to (2), the following condition:
$G^{-}_{-1/2}\Phi^{l}_{q}=0.$ (7)
These states are BPS, so the additional constraints they obey require $\Delta^{NS}=\frac{|Q^{NS}|}{2}$ Lerche:1989uy .
Recall that in minimal model the chiral-primary and anti-chiral-primary states
@@
appear when $q=\pm l$.
In the models of $N=(2,2)$ SCFT, the primary fields are composed of the appropriately paired holomorphic and anti-holomorphic primary states. In the modular bootstrap approach, the particular way of pairing can be read off from the partition function. As it was found by D.Gepner Gepner:1988 the most general modular invariant partition function of the minimal model takes the form
$Z(\tau)=\frac{1}{2}\sum_{l,\bar{l}=0}^{k}\sum_{\begin{subarray}{c}q,\bar{q}=-(k+1);\\ l+q+s=0,\ \bar{l}+\bar{q}+\bar{s}=0\ \text{mod}\ 2\end{subarray}}^{k+2}\sum_{\begin{subarray}{c}s,\bar{s}=0\end{subarray}}^{3}N_{l,\bar{l}}M_{q,\bar{q}}L_{s,\bar{s}}\chi_{q,s}^{l}(\tau)\bar{\chi}_{\bar{q},\bar{s}}^{\bar{l}}(\bar{\tau}),$ (9)
where $\chi_{q,s}^{l}$, $s=0,1,2,3$ are the splited characters of the highest weight representation $\mathcal{H}^{l}_{q}$ of $N=2$ Virasoro superalgebra, which are calculated over the subspaces with a fixed fermion number. More precisely, $s=0,1$ give the characters calculated over the subspaces with even number of fermionic generators of $G^{\pm}$ in $NS$ and $R$ sectors correspondingly, while $s=2,3$ give the characters calculated over the subspaces with odd number of fermionic generators of $G^{\pm}$ in $NS$ and $R$ sectors. The matrices $N_{l,\bar{l}},M_{q,\bar{q}},L_{s,\bar{s}}$ provide modular invariance of the partition function (see also Gepner:Parafermion , Gepner:Lectures ).
@@
which includes the simplest diagonal invariant of type $\textbf{A}_{k+1}$ for all integer levels $k\geq 1$. For even $k$, two series of off-diagonal invariants of type D arise: $\textbf{D}_{2j+2}$ for $k=4j$, and $\textbf{D}_{2j+1}$ for k=4j2k=4j-2. Besides these infinite series, there are also three exceptional cases: $\textbf{E}_{6},\textbf{E}_{7},$ and $\textbf{E}_{8}$ for $k=10,16,$ and 28 respectively. The alternative proof of the classification (10) can be found in Gannon:1999cp .
In the very nice paper Gannon:1996hp the more general list of modular invariants for the $N=(2,2)$ superconformal minimal models has been found. But in this paper we restrict ourself to the models whose partition functions are given by (9), where the matrix $N_{l,\bar{l}}$ is given by (10) and the matrices $M$, $L$ are identity matrices.
It is worth noting that the expression for the partition function in (9) is, in a sense, formal. In it the parameter $q$ goes through all possible values $q=-(k+1),\dots,(k+2)$ and depends on $l$ through the parity condition: l+q=0mod 2l+q=0\ $\text${mod}\ 2 in $NS$ sector, and l+q=1mod 2l+q=1\ $\text${mod}\ 2 in $R$ sector. However, the minimal representations of the $N=2$ Super-Virasoro algebra are defined only for $q=-l,-l+2,\dots,l$. Consequently, the corresponding highest vector $\Phi^{l}_{q}$ of the representation $\mathcal{H}^{l}_{q,s}$ with $|q|>l$ does not belong to the Kac table of the minimal model $M_{k}$ and their conformal dimension and $U(1)$ charge can not be computed using (4). Despite this, the symbolic form of (9) facilitates the proof of modular invariance, since the indices $(l,\bar{l})$, $(q,\bar{q})$, and $(s,\bar{s})$ in the product $N_{l,\bar{l}}M_{q,\bar{q}}L_{s,\bar{s}}$ decouple under modular transformations Gepner:1988 .
To clarify the interpretation of the remaining terms in the partition function, we can rely on the isomorphisms of $N=2$ Virasoro superalgebra representations Gepner:1987vz , Gepner:Lectures , Feigin , established rigorously in FSST as the isomorphisms of twisted representations:
@@
These isomorphisms allow us to clarify the meaning of the terms appearing in (9) that correspond to the states $\Phi^{l}_{q}$ with $|q|>l$. They should be understood via the identifications
$\Phi_{q}^{l}\simeq\Phi_{q+(k+2)}^{k-l},\quad\Phi_{q}^{l}\simeq\Phi_{q+2(k+2)}^{l}.$ (12)
Below in our article, especially in section 5 with examples, the states $\Phi^{l}_{q}$ with $|q|>l$ may arise. They should be understood as being identified with $\Phi^{k-l}_{q\pm(k+2)}$. For now, the identifications (11), (12) are more of an enumerative nature. Below, in section 3, we will explain the exact meaning of these identifications.
Isomorphisms of representations (11) provide equality of corresponding characters
$\chi_{q,s}^{l}(\tau)=\chi_{q+(k+2),s+2}^{k-l}(\tau),\quad\chi_{q}^{l}(\tau)=\chi_{q+2(k+2),s+4}^{l}(\tau).$ (13)
Using these it is easy to explain the coefficient $\frac{1}{2}$ in the partition function (9). The point is that the parameters $(l,q,s)$ and $(\bar{l},\bar{q},\bar{s})$ run over all possible values, so there are terms with characters $\chi_{q,s}^{l}(\tau)$ and $\chi_{q+(k+2),s+2}^{k-l}(\tau)$. Due to their equality, each term appears twice. However, this is not related to multiplicity, but arises due to the symbolic notation of the partition function (9), as we explained above. The only exception is the D series for $k=4j$. In it the multiplicity of characters with $l=\bar{l}=\frac{k}{2}$ is 2. This amazing fact has been discussed, for example, in the paper Fuchs:1991vu .
@@
Of course the holomorphic factorization we depicted in (14) should not be taken literally. The summation of quantum group indices, which ensures mutual locality of fields within correlation functions, is certainly implied and will be implied in subsequent similar expressions. But these details are not very important for the discussion in the paper. So, in what follows we will use rather symbolic expressions like (14) instead of explicit formulas.
Taking into account the identification (12) we can assume that in $NS$ sector the charge $q$ runs:
$q=-(k+1),\dots,(k+2),\quad q+l=0\ (\text{mod}\ 2),$ (15)
and in $R$ sector:
$q=-(k+1),\dots,(k+2),\quad q+l=1\ (\text{mod}\ 2).$ (16)
@@ 2.1 Example: spectrum of $\textbf{E}_{7}$ minimal model
Let us show how one can find the spectrum of the minimal model, defined by the modular invariant (9) using the example of the invariant $\textbf{E}_{7}$ with $k=16$. The partition function of the corresponding $\widehat{sl}(2)$ WZW model is
$Z_{E_{7}}=\left|\chi_{0}+\chi_{16}\right|^{2}+\left|\chi_{4}+\chi_{12}\right|^{2}+\left|\chi_{6}+\chi_{10}\right|^{2}+\left|\chi_{8}\right|^{2}+\left(\chi_{2}+\chi_{14}\right)\bar{\chi}_{8}+\chi_{8}\left(\bar{\chi}_{2}+\bar{\chi}_{14}\right).$ (17)
So, the matrix $N_{l,\bar{l}}$ of the Minimal model can be read off from here. Since the matrices $M$, $L$ are identity matrices, we can easily write out the list of $NS$ primary fields of the Minimal model $\textbf{E}_{7}$ taking into account the identifications (12). This set is shown in table 1.
The first column contains the term from the partition function, which determines the corresponding matrix elements $N_{l,\bar{l}}$ and corresponds to the fields in the second column. Since these are fields from the $NS$ sector, $q+l$ is even for all fields in the table. Note that some fields have left and right states with different charges: $\bar{q}=q\pm(k+2)$. It arises from the identification rules (12) when we bring the labels $q,\bar{q}$ in to the domain $|q|\leq l,|\bar{q}|\leq\bar{l}$. It is also straightforward to list the primary fields of the $R$-sector as well.
@@ 3 Orbifolds of composite models
The fields above, of course, do not exhaust the entire list of primary fields of the composite model. However, they are especially important for us since all other fields in the composite model can be obtained from them by applying the operators of the $N=(2,2)$ Virasoro superalgebras of $M_{k_{i}}$ models.
@@ 3.1 Group of symmetry of the composite model
$G_{\text{tot}}=\prod_{i=1}^{r}G_{i}=\left\{\prod_{i}\hat{g}_{i}^{w_{i}},w_{i}\in\mathbb{Z},\hat{g}_{i}=\exp\left(i2\pi J_{(i),0}\right)\right\},$ (21)
where $G_{i}$ is the discrete symmetry group of the individual model $M_{k_{i}}$. The $G_{\text{tot}}$ depends on the modular invariant (10) in each copy of $M_{k_{i}}$.
In the $\textbf{D}_{2j+2}$ type series for $k=4j$, as well as in $\textbf{E}_{7}$ and $\textbf{E}_{8}$ the parameters $l_{i},\bar{l}_{i}$ run through only even values. Then the charges $q_{i}$ of primary state are also only even. It follows that for the corresponding element $\hat{g}_{i}^{w_{i}}\in G_{i}$, the exponent $w_{i}$ is defined modulo $(k_{i}+2)/2$, then $G_{i}\simeq\mathbb{Z}_{(k_{i}+2)/2}$. For modular invariants of type $\textbf{A}_{k_{i}+1}$ and $\textbf{D}_{2j+1}$ with $k_{i}=4j-2$ as well as for the exceptional $\textbf{E}_{6}$, there are no restrictions on the parity of the charges. Then the corresponding group is $G_{i}\simeq\mathbb{Z}_{k_{i}+2}$. The case of diagonal invariants BBP has a group $G_{\text{tot}}=\prod_{i=1}^{r}\mathbb{Z}_{k_{i}+2}$.
Below in the text, it will be useful to use additive version of the group $G_{tot}$ labeling its elements by the vectors $\vec{w}=(w_{1},...,w_{r})$ and regard the elements of the group $w_{i}\in\mathbb{Z}_{(k_{i}+2)/2}$ as even elements in $\mathbb{Z}_{k_{i}+2}$. We will explain this restriction in more detail in section 4.
The admissible group is defined as any subgroup of $G_{\text{tot}}$, satisfying
$G_{\text{adm}}=\left\{\vec{w}=(w_{1},\dots,w_{r})\bigg|\quad w_{i}\in G_{i},\quad\sum_{i}^{r}\frac{w_{i}}{k_{i}+2}\in\mathbb{Z}\right\}\subset G_{tot}.$ (22)
As will be seen (see section 5), the constraint
@@ 3.2 Spectral Flow twisting and mutual locality of fields
In order to implement the first step, we use the spectral flow realization of primary states BP ; BBP ; IIAB . It goes as follows.
@@
For values $t\in\mathbb{Z}+\frac{1}{2}$ the spectral flow maps $NS$ and $R$ sectors into each other, and for $t\in\mathbb{Z}$ it acts inside $NS$ or $R$ sectors.
The extension of the spectral flow automorphism above to the $N=2$ Virasoro superalgebra representations Lerche:1989uy generates so-called twisted representations FSST . In particular, the initial highest weight vector $\Phi^{l}_{q}$ becomes the twisted ones after the spectral flow operator $U^{t}$ action. At the same time, exactly the characters of these twisted representations appear in Gepner’s modular invariant partition function of superstring compactification. It raises the question of whether we can find the standard highest weight vector among the states of the twisted representations?
@@
gives spectral flow realization of the $NS$ primary state $\Phi^{l}_{q}$ of the individual minimal model, where q=l2tq=l-2t. In other words, the expression (25) gives the construction of the primary state $\Phi^{l}_{q}$ as a specific descendant in the chiral-primary representation with the chiral-primary state $\Phi^{l}_{c}$, which is twisted by $U^{t}$, where $0\leq t\leq l$.
For the purposes of the orbifold construction, we need to extend this formula. Namely, the state
@@
gives the spectral flow realization of the $NS$ primary state $\Phi^{\tilde{l}}_{\tilde{q}}$, where $\tilde{l}=k-l$, $\tilde{q}=k+2+l-2t$. The expression (26) gives the construction of the primary state $\Phi^{\tilde{l}}_{\tilde{q}}$ as a specific descendant in the chiral-primary representation with the chiral-primary state $\Phi^{l}_{c}$, which is twisted by $U^{t}$, where $l+1\leq t\leq k+1$.
It is clear that the $R$ primary state is given by $U^{\frac{1}{2}}$ twisting of the $NS$ primary fields above. The validity of formulas (25), (26) follows from the structure of singular vectors of unitary, integrable representations of the $N=2$ super-Virasoro algebra FSST , Feigin . Below we generalize these formulas to construct the primary fields in the orbifold model.
Notice that for $t=k+2$ the state
$V^{l}_{k+2}=U(UG^{-}_{-\frac{1}{2}})^{k-l}U(UG^{-}_{-\frac{1}{2}})^{l}\Phi^{l}_{c}$ (27)
gives the spectral flow realization of the primary state which can be identified with the original chiral primary one $\Phi^{l}_{c}$. From this perspective, the action of the spectral flow operator on the representations has a ”periodicity” condition
$U^{k+2}\simeq 1.$ (28)
As it is clear from (27) this relation should not be interpreted literally but as an isomorphism between representations generated by highest vectors (25)-(26) at $t=0$ and $t=k+2$. Consequently, the parameter $t$ is defined modulo $k+2$.
The previously mentioned identification of states (12), should also be understood within the framework of the relations (25),(26). Indeed, while (25) provides the spectral flow realization of the primary state $\Phi^{l}_{q}$, where q=l2tq=l-2t and hence, $|q|\leq l$, one can consider the expression (26) as giving spectral flow realization of some primary state $\Phi^{l}_{q}$, where again q=l2tq=l-2t, but this time $l\leq|q|$. From the other hand, this state is nothing else but the primary state $\Phi^{k-l}_{q+k+2}$, where $|q+k+2|\leq k-l$. It yields the first identification from (12). This implies also isomorphisms of representations (11). Given (27), the second identification from (12) should be understood similarly.
We also note that it is possible to extend the domain of the spectral flow parameter $t$ to the full set of integers and obtain an infinite set of realizations of the primary states by appropriately generalizing the expressions (26). This amazing fact has already been implied in FST and is certainly worth being considered further. However, we leave this issue beyond the scope of the present paper.
In terms of spectral flow realization of the primary states, the primary field $\Phi_{q}^{l}(z)\bar{\Phi}^{\bar{l}}_{q}(\bar{z})$ from the minimal model $M_{k}$ with explicit ADE invariant can be realized as follows
@@
Now we use the spectral flow realizations (25), (26), (29) to build twisted states from the states of the composite model for each element of the admissible group $\vec{w}\in G_{\text{adm}}$.
In $NS$ sector the twisted primary field is given by
@@
$\bar{V}_{\vec{\bar{t}}}^{\vec{\bar{l}}}(\bar{z})=\prod_{i}^{r}\bar{V}_{\bar{t}_{i}}^{\bar{l}_{i}}(\bar{z}),\quad\bar{V}_{\bar{t}_{i}}^{\bar{l}_{i}}(z)=\left(UG_{-\frac{1}{2}}^{-}\right)_{i}^{\bar{t}_{i}}\bar{\Phi}^{\bar{l}_{i}}_{c}(z),\quad 0\leq\bar{t}_{i}\leq\bar{l}_{i}$ (31)
@@
Note that for $\vec{w}=0$, the expressions (30)-(33) give the spectral flow construction of primary fields (20) from $NS$ sector of the original composite model, where $\vec{q}=\vec{l}-2\vec{t}=\vec{\bar{l}}-2\vec{\bar{t}}$.
@@
Then the primary states of the minimal model factors of the composite model, as is well known, can be represented as ZF
@@
where the $\hat{\Phi}_{q_{i}}^{l_{i}}(z)$ is uncharged with respect to current $J_{i}(z)$. The complete primary field of the model $M_{k_{i}}$ is an appropriate product of the factor above and an antiholomorphic factor that can be represented similarly (as we have already mentioned, this product must include quantum group indices summation).
@@
$\exp\left[i\sum_{i}\left(\sqrt{\frac{k_{i}}{k_{i}+2}}w_{1i}+\frac{q_{1i}}{\sqrt{k_{i}\left(k_{i}+2\right)}}\right)\phi_{i}(z)\right]\times\\ \times\exp\left[i\sum_{j}\left(\sqrt{\frac{k_{j}}{k_{j}+2}}w_{2j}+\frac{q_{2j}}{\sqrt{k_{j}\left(k_{j}+2\right)}}\right)\phi_{j}(0)\right],$ (38)
where $q_{1,2i}=l_{1,2i}-2t_{1,2i}$. Applying the well-known result for exponents of free fields:
@@
There is another way to get this equation which doesn’t use the representation (37). To this end one can consider the operator product expansion of the fields (34) and demand the corresponding monodromy vanishes (see Appendix in IIAB ).
The equation (40) is completely analogous to the equation obtained for composite models of type A in BBP ; Parkhomenko:2022kju . Thus, the off-diagonality of the fields (30) with respect to the indices $\vec{l},\vec{\bar{l}}$, arising for D and E invariants, affects the mutual locality condition only through the allowed set of vectors $\vec{q}$ which appear in the twisted sector $\vec{w}$.
@@
Thus, by solving the mutual locality equations (40) with respect to $\vec{q}$ in each twisted sector $\vec{w}$ and taking into account descendants, we can explicitly determine all fields of the orbifold model. It is easy to see that in this way we have obtained a set of (quasi-)local fields that also satisfy the other axioms of the conformal bootstrap. In other words, we obtain a model of a $N=(2,2)$ superconformal field theory.